OpenStudy (pandaecat):

Simplify square root of 2 multiplied by the cube root of 2.

OpenStudy (pandaecat):

Can someone explain to me please???

OpenStudy (pandaecat):

Do you solve it like this: square root of 2 multiplied by the cube root of 2 equals2^(1/2)+2^(1/3) then you just solve 1/2 + 1/3 which equals 5/6??? So the answer is 2 and 5/6???

Nnesha (nnesha):

convert root to exponent form and then use the exponent rule when you multiply the same bases you should add their exponents

Nnesha (nnesha):

2 and 5/6 how would you write it in math form

OpenStudy (pandaecat):

Math form???

Nnesha (nnesha):

so it is like 2 and 5/6 like 2 times 5/6 what is 5/6 exponent ?

OpenStudy (pandaecat):

I don't really get it >.<

OpenStudy (pandaecat):

 that's what I meant by the answer

Nnesha (nnesha):

\[\huge\rm 2^\frac{ 1 }{ 3 } \times 2^\frac{ 1 }{ 2 } =2^{\frac{ 1 }{ 2 }+\frac{ 1 }{ 3 }}\] so 5/6 is exponent of 2

Nnesha (nnesha):

\[\huge\rm 2^\frac{ 5 }{ 6 }\] now you can convert 5/6 exponent to rot

Nnesha (nnesha):

root*

OpenStudy (pandaecat):

Oh I see!!!

Nnesha (nnesha):

\[\huge\rm \sqrt[m]{x^n} = x^\frac{ n }{ m }\]